[Hallc_running] [New Logentry] Hall C RC Update: Wed., Jan. 12, 2022

gaskelld at jlab.org gaskelld at jlab.org
Wed Jan 12 08:45:01 EST 2022

Logentry Text:
The target was successfully cooled down by yesterday afternoon, although some accelerator-related issues delayed beam delivery into owl shift.  On owl shift, there were two faults of the SHMS Q1 lead-flow controller - a Controlled Access has been requested to work on this today.  The beam-target centering was completed on owl shift, although a pretty significant change in the horizontal beam position was required. 

After the Controlled Access, we'll start production on the Q2=3.85, W=3.07, low epsilon setting.

There will be an RC meeting at 10 am this morning:


Zoom link:



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3970459
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220112/b13c8569/attachment.html>

More information about the Hallc_running mailing list